OpenStudy (anonymous):

(-6,3); 8x=3y+5 Express answer in the form y=mx+b

OpenStudy (anonymous):

\[8x - 5= 3y + 5 - 5\]\[8x - 5= 3y \] \[(8x - 5)/3= 3y/3 \] \[(8x/3) - (5/3)= y \]